# SoK: Cybersecurity Assessment of Humanoid Ecosystem

### v1

@@ Abstract
Humanoids are progressing toward practical deployment across healthcare, industrial, defense, and service sectors. While typically considered cyber-physical systems (CPSs), their dependence on traditional networked software stacks (e.g., Linux operating systems), robot operating system (ROS) middleware, and over-the-air update channels, creates a distinct security profile that exposes them to vulnerabilities conventional CPS models do not fully address. Prior studies have mainly examined specific threats, such as LiDAR spoofing or adversarial machine learning (AML). This narrow focus overlooks how an attack targeting one component can cascade harm throughout the robot’s interconnected systems. We address this gap through a systematization of knowledge (SoK) that takes a comprehensive approach, consolidating fragmented research from robotics, CPS, and network security domains. We introduce a seven-layer security model for humanoid robots, organizing 39 known attacks and 35 defenses across the humanoid ecosystem—from hardware to human–robot interaction. Building on this security model, we develop a quantitative 39×35 attack–defense matrix with risk-weighted scoring, validated through Monte Carlo analysis. We demonstrate our method by evaluating three real-world robots: Pepper, G1 EDU, and Digit. The scoring analysis revealed varying security maturity levels, with scores ranging from 39.9% to 79.5% across the platforms. This work introduces a structured, evidence-based assessment method that enables systematic security evaluation, supports cross-platform benchmarking, and guides prioritization of security investments in humanoid robotics.
@@ 1 Introduction
Humanoids. Humanoids are being deployed in sectors such as healthcare, industrial automation, defense, and service sectors by replicating human motion, cognition, and interaction [100, 50, 120]. Powered by AI, these robots are designed for real-time, high-stakes tasks such as surgical assistance, elder care, and public interaction. They also enable close collaboration in dynamic human environments [124]. However, the very capabilities that make humanoids so promising also expose them to new cybersecurity risks. Before these systems are widely deployed, it is critical to establish a rigorous method for assessing their security. To address this challenge, we present a comprehensive systematization of knowledge (SoK) that introduces a seven-layer security architecture model specifically designed for humanoid robots. With help of this model, we organize 39 documented attacks and 35 defense mechanisms across the humanoid technology stack, from hardware components to human-robot interaction layers. The purpose of this security model is to enable systematic vulnerability assessment, support cross-platform security benchmarking, and guide evidence-based security investment decisions for humanoid deployments.
@@
A unique and layered threat landscape. The security challenges faced by humanoids are fundamentally different from those of classical CPSs or traditional IT systems. Humanoids must navigate dynamic, unpredictable settings with constantly shifting variables like human movement and environmental obstacles [104]. Their integration into human-centered environments introduces a unique threat landscape, where cyber attacks on humanoid systems can jeopardize user safety, compromise privacy, and disrupt operational reliability [36, 117]. These security breaches may lead to direct and immediate physical consequences.
@@
This threat landscape is exacerbated by the deeply interconnected nature of humanoid architecture, where multi-layered subsystems create cross-layer dependencies and an expansive attack surface [73, 58, 56, 114, 22]. Unlike traditional CPSs, humanoids integrate numerous attack-prone subsystems: AI accelerators, sensor arrays, communication middleware, and decision-making algorithms—each with distinct vulnerability profiles that can cascade failures across the entire system. At the hardware level, humanoids face firmware tampering and sensor spoofing, with their AI accelerators (e.g., Jetson Orin, Neural Processing Units (NPUs)) being potential targets for exploits that compromise system integrity [92]. In the decision-making area, their reliance on deep learning for navigation and decision-making makes them vulnerable to adversarial and data poisoning attacks, which can trigger unpredictable or dangerous behaviors [53, 68]. Finally, their communication and middleware infrastructure, often based on ROS 2, provides openings for man-in-the-middle attacks, unauthorized control hijacking, and real-time data manipulation [19, 34]. Mitigating these varied, layered threats requires a comprehensive and specialized security approach.
@@
The examples highlights the fact that modern humanoid architectures are susceptible to the security vulnerabilities discussed earlier. As these robots become increasingly autonomous and integrated into our daily lives, a systematic approach for identifying and mitigating these risks is essential for ensuring humanoids’ safe and trustworthy operation.
@@ 2 Related Work and Research Gaps
Humanoid security research spans robotics, AML, CPSs, and human-computer interaction domains. However, existing research remains fragmented across these disciplines, lacking unified architectural models and systematic analyses of cross-layer vulnerability propagation. Most prior work has focused on isolated threats without considering how attacks targeting one subsystem can cascade throughout the interconnected humanoid architecture. Our SoK addresses this fragmentation by conducting systematic literature analysis to answer three key research questions: (1) How do existing security threats cluster across humanoid architectural boundaries? (2) What gaps exist in current attack-defense coverage? (3) Can architectural patterns from the literature help in systematic security assessment approaches for humanoid platforms? Our analysis synthesizes 100 high-confidence studies to reveal consistent architectural patterns that inform the development of a structured security model specifically designed for humanoid systems.
Research methodology and paper selection. We conducted a PRISMA-compliant systematic review targeting humanoid security research across four major academic databases (IEEE Xplore, ACM Digital Library, SpringerLink, and arXiv) using the search keywords: ("humanoid robot" OR "social robot" OR "service robot") AND ("security" OR "cybersecurity" OR "adversarial" OR "attack"), spanning publications from 2015-2024. Our inclusion criteria required: (1) explicit security content with concrete threat analysis, (2) embodied robotics relevance beyond theoretical security models, and (3) peer-reviewed publications or high-quality preprints. Studies were excluded if they focused solely on theoretical security concepts without application to robotics, or if they pertained to industrial automation lacking humanoid characteristics. The initial search yielded 547 papers. After title/abstract screening, 150 papers underwent full-text analysis, resulting in 100 high-confidence studies that form our research corpus. Through systematic content analysis, we identified 39 distinct attack vectors and 35 defense mechanisms, which clustered around seven functional architectural layers. This corpus demonstrates significant temporal evolution: 68% of relevant work was published after 2020, indicating rapid field emergence and the need for systematic organization of this growing knowledge base.
@@
While demonstrating practical exploitability, these platform-specific studies lack the architectural abstraction necessary for systematic cross-platform analysis and fail to model how vulnerabilities propagate across the layered subsystems that comprise integrated humanoid architectures. We observed that existing approaches share a common limitation: they analyze security within traditional domain boundaries rather than recognizing the natural architectural layers of humanoid system design. This gap motivates our development of a layered architectural model that organizes security analysis around the functional boundaries inherent to embodied autonomous systems.
2.1 Current State: The Fragmented Security Landscape
Our systematic analysis of 100 high-confidence papers reveals a research landscape characterized by domain-specific depth but limited cross-domain integration. While individual security domains have developed sophisticated attack-defense frameworks, the lack of unified architectural thinking has created critical blind spots in understanding system-wide vulnerabilities.
AML in Robotics. The intersection of AML and robotics has emerged as a active research area, though coverage remains narrow. Recent work demonstrates sophisticated attacks against core AI components: visual simultaneous localization and mapping (SLAM) systems have been compromised through carefully crafted perturbations that cause localization failures in navigation tasks, while reinforcement learning policies have been manipulated to produce suboptimal or dangerous actions during critical decision-making phases [15, 13]. Sequential perturbation attacks against quadrupedal locomotion systems have shown how adversaries can induce gait instabilities that compromise mobility, while supply chain attacks have been demonstrated against vision-language models used for robotic manipulation tasks [86, 110]. The study of these work indicated that while attack methodologies are well-documented across the surveyed papers, corresponding defense mechanisms appear in a few, creating a significant vulnerability gap. Moreover, these studies focus exclusively on isolated AI components without considering how AML attacks might propagate across integrated robotic architectures.
Middleware and Communication Security. Within the reviewed papers, ROS and data distribution service (DDS) security were found to be most studied topics, with approximately one-third of the surveyed publications addressing both attack vectors and corresponding defense mechanisms [19, 78, 3]. Researchers have found that ROS and DDS are vulnerable to attacks like unauthorized command injection [93, 122], and some defenses such as intrusion and anomaly detection [80] have been proposed. However, the current research treats middleware as an isolated layer, failing to model how communication vulnerabilities interact with sensing, planning, and actuation subsystems in real-world deployments.
@@ 2.2 Recurring Architectural Themes
Our systematic analysis uncovered a distinct pattern: despite the apparent diversity of humanoid platforms and attack methodologies examined across papers, security concerns consistently clustered around seven distinct functional boundaries. During our content analysis, threats that initially appeared unrelated; consistently aligned with the same architectural layers. For example, sensor-spoofing attacks, whether targeting LiDAR systems [88, 14] or cameras [98, 24], invariably exploited vulnerabilities at the sensory interface between physical hardware and digital processing. These attacks succeeded by manipulating the boundary where analog sensor data transforms into digital representations used by higher-level algorithms. Similarly, AML attacks consistently targeted high-level reasoning components, exploiting the interface between data processing and decision-making logic [15, 13, 86]. Physical components, upon which all other system functions depend, can be compromised by hardware attacks. Most significantly, the observed patterns reflect real architectural boundaries in humanoid robots—not merely the areas researchers emphasized. Unlike traditional IT systems with arbitrary software boundaries, humanoids have natural layers that reflect their physical and digital components. These layers create consistent attack points where different subsystems connect, making them common targets across different platforms.
@@ 2.3 Critical Gaps
Our literature analysis shows that research on robotics security is highly fragmented. The vast majority of studies performed have focused on specific attack vectors or individual subsystems; only a few of the surveyed papers examined attacks spanning more than one subsystem or architectural layer. As a result, prior work often fails to capture realistic adversarial dynamics, where a local fault can propagate across the sensing, the planning, and the actuation layers.
We observe an asymmetric defense landscape. Human–robot interface (HRI) vulnerabilities are well documented [122, 119], but practical, validated mitigations are scarce. By contrast, several sensor-level attacks (e.g., LiDAR spoofing, IMU drift) [88, 14] already have concrete countermeasures (filtering, fusion checks, shielding, freshness tests) [109, 116]. The result is an imbalanced posture: upper layers that shape user intent are weakly protected, while lower layers have spot-fixes. Modern robotic systems need more measures like integrated defenses that work together across all system layers. This means binding human-robot interaction policies to authenticated, session-verified middleware; requiring sensor validation before any planning occurs; and implementing hard limits at the control level with automated recovery capabilities. The proposed model makes this integration measurable by identifying the specific layers and interfaces that must connect, while RISK-MAP quantifies whether a platform actually implements these connections and calculates the residual risk when any link fails.
@@ 2.4 Contributions of This Paper
    Novel seven layered framework: Introduction of seven-layer security model tailored to humanoids that organizes 39 attack vectors and 35 defenses across natural system boundaries identified through our literature analysis.
@@
    Quantitative assessment methodology: Introduction of a risk-weighted scoring system with statistical validation that enables systematic security evaluation and cross-platform benchmarking.
@@ 3.1 Architectural Analysis
The network-CPS duality. Our analysis uncovered that humanoids exist at the intersection of two established computational paradigms. Like networked systems, they use layered abstractions with defined interfaces, protocol stacks, and modular data flows. Like CPSs, they couple computation to physical processes under real-time and safety constraints. This duality implies that an effective security architecture must treat interfaces as enforcement points for identity, authorization, integrity, confidentiality, and auditability and practically, use each interface to enforce who can communicate, what can be send, whether it is within safe bounds, and recorded for audit.
@@
Therefore, we developed a seven-layer security architecture tailored to humanoids by refining the OSI model, in an iterative process. This process involved following steps:
@@
    Defense Integration: Examining how the 35 defense mechanisms in the literature review can be systematically organized in the layered framework.
@@
3.2 7-Layer Humanoid Security Model
@@ 4.1 Holistic Humanoid Threat Landscape
Real-time constraints make latency itself an attack vector. In enterprise IT, a few seconds of delay in anomaly detection may be tolerable; in robotics, locomotion and manipulation often run faster than 10 milliseconds, so the same delay can cause physical collapse, hardware damage, or injury. Adversaries can exploit this narrow operational window, for example, LiDAR spoofing (SP-A1) can destabilize motion within a control cycle, before a detector reacts. Security mechanisms that do not operate within these deadlines provide no prevention—they enable only forensic analysis.

### v2

@@ Abstract
Humanoids are progressing toward practical deployment across healthcare, industrial, defense, and service sectors. While typically considered cyber-physical systems (CPSs), their dependence on traditional networked software stacks (e.g., Linux operating systems), robot operating system (ROS) middleware, and over-the-air update channels, creates a distinct security profile exposing them to vulnerabilities that conventional CPS models do not fully capture. Prior studies have mainly examined specific threats, such as LiDAR spoofing or adversarial machine learning (AML). This narrow focus overlooks how an attack targeting one component can cascade harm throughout the robot’s interconnected systems. We address this gap through a systematization of knowledge (SoK) that takes a comprehensive approach, consolidating fragmented research from robotics, CPS, and network security domains. We introduce a seven-layer security model for humanoid robots, organizing 39 known attacks and 35 defenses across the humanoid ecosystem-from hardware to human-robot interaction. Building on this security model, we develop a quantitative 39×35 attack–defense matrix with risk-weighted scoring, validated through Monte Carlo analysis. We demonstrate our method by evaluating three real-world robots: Pepper, G1 EDU, and Digit. The scoring analysis revealed varying security maturity levels ranging from 39.9% to 79.5% across the platforms. This work introduces a structured, evidence-based assessment method that enables systematic security evaluation, supports cross-platform benchmarking, and guides prioritization of security investments in humanoid robotics.
@@ 1 Introduction
Humanoids. Humanoids are being deployed in sectors such as healthcare, industrial automation, defense, and service sectors by replicating human motion, cognition, and interaction [97, 53, 28]. Powered by AI, these robots are designed for real-time, high-stakes tasks such as surgical assistance, elder care, and public interaction. They also enable close collaboration in dynamic human environments [38]. However, the very capabilities that make humanoids so promising also expose them to new cybersecurity risks. Before these systems are widely deployed, it is critical to establish a rigorous method for assessing their security. To address this challenge, we present a comprehensive systematization of knowledge (SoK) that introduces a seven-layer security architecture model specifically designed for humanoid robots. With the help of this model, we organize 39 documented attacks and 35 defense mechanisms across the humanoid technology stack, from hardware components to human-robot interaction layers. The purpose of this security model is to enable systematic vulnerability assessment, support cross-platform security benchmarking, and guide evidence-based security investment decisions for humanoid deployments.
@@
A unique and layered threat landscape. The security challenges faced by humanoids are fundamentally different from those of classical CPSs or traditional IT systems. Humanoids are perimeter-less assets: they work outside secured facilities and can be approached directly by humans in hospitals, factories, and public venues. This reality elevates the importance of Physical and Social interaction and reshapes threat priorities beyond traditional IT/CPS assumptions. Humanoids must navigate dynamic, unpredictable settings with constantly shifting variables like human movement and environmental obstacles [105]. Their integration into human-centered environments introduces a unique threat landscape, where cyber attacks on humanoid systems can jeopardize user safety, compromise privacy, and disrupt operational reliability [40, 118]. These security breaches may lead to direct and immediate physical consequences.
@@
This threat landscape is exacerbated by the deeply interconnected nature of humanoid architecture, where multi-layered subsystems create cross-layer dependencies and an expansive attack surface [78, 64, 62, 115, 22]. Unlike traditional CPSs, humanoids integrate numerous attack-prone subsystems: AI accelerators, sensor arrays, middleware, and decision-making algorithms-each with distinct vulnerability profiles that can cascade failures across the entire system. At the hardware level, humanoids face firmware tampering and sensor spoofing, with their AI accelerators (e.g., Jetson Orin, Neural Processing Units (NPUs)) being potential targets for exploits that compromise system integrity [95]. In the decision-making area, their reliance on deep learning for navigation and decision-making makes them vulnerable to adversarial and data poisoning attacks, which can trigger unpredictable or dangerous behaviors [58, 73]. Finally, their communication and middleware infrastructure, often based on ROS 2, provides openings for man-in-the-middle attacks, unauthorized control hijacking, and real-time data manipulation [19, 37]. Mitigating these varied, layered threats requires a comprehensive and specialized security approach.
@@
The examples highlight the fact that modern humanoid architectures are susceptible to the security vulnerabilities discussed earlier. As these robots become increasingly autonomous and integrated into our daily lives, a systematic approach for identifying and mitigating these risks is essential for ensuring humanoids’ safe and trustworthy operation.
@@ 2 Related Work and Research Gaps
Humanoid security research spans robotics, AML, CPSs, and human-computer interaction domains. However, existing research remains fragmented across these disciplines, lacking unified architectural models and systematic analyses of cross-layer vulnerability propagation. Most prior work has focused on isolated threats without considering how attacks targeting one subsystem can cascade throughout the interconnected humanoid architecture. Our SoK addresses this fragmentation by conducting systematic literature analysis to answer three key research questions: (1) What are the primary security threats facing humanoid robots and how do they relate to its system architecture? (2) What gaps exist in current attack-defense coverage? (3) Can architectural patterns from the literature help in systematic security assessment approaches for humanoid platforms? Our analysis synthesizes 89 high-confidence studies to reveal consistent architectural patterns that inform the development of a structured security model specifically designed for humanoid systems.
Research methodology and paper selection. We conducted a PRISMA-compliant systematic review targeting humanoid security research across four major academic databases (IEEE Xplore, ACM Digital Library, SpringerLink, and arXiv) using the search keywords: ("humanoid robot" OR "social robot" OR "service robot") AND ("security" OR "cybersecurity" OR "adversarial" OR "attack"), spanning publications from 2015-2025. Our inclusion criteria required: (1) explicitly address security with concrete threat analyses, (2) embodied robotics relevance beyond theoretical security models, and (3) peer-reviewed publications or high-quality preprints. Studies were excluded if they focused solely on theoretical security concepts without application to robotics, or if they pertained to industrial automation lacking humanoid characteristics. The initial search yielded 530 papers. After title/abstract screening, 150 papers underwent full-text analysis, resulting in 89 high-confidence studies that form our research corpus. Through systematic content analysis, we identified 39 distinct attack vectors and 35 defense mechanisms, which clustered around seven functional architectural layers. This corpus shows rapid growth, with 68% published after 2020, indicating rapid field emergence and the need for systematic organization of this growing knowledge base.
@@
While demonstrating practical exploitability, these platform-specific studies lack the architectural abstraction necessary for systematic cross-platform analysis and fail to model how vulnerabilities propagate across the layered subsystems that comprise integrated humanoid architectures. Our review confirms that most studies examine isolated threats (sensor spoofing, adversarial ML) rather than modeling how they cascade across architectures. This gap motivates our development of a layered architectural model that organizes security analysis around the functional boundaries inherent to embodied autonomous systems.
2.1 Current State: Siloed Security Approaches
Our systematic analysis of 89 high-confidence papers reveals a research landscape characterized by domain-specific depth but limited cross-domain integration. While individual security domains have developed sophisticated attack-defense frameworks; they remain bounded by traditional domains, limiting their ability to capture cross-layer propagation in understanding system-wide vulnerabilities.
AML in Robotics. The intersection of AML and robotics has emerged as an active research area, though coverage remains narrow. Recent work demonstrates sophisticated attacks against core AI components: visual simultaneous localization and mapping (SLAM) systems have been compromised through carefully crafted perturbations that cause localization failures in navigation tasks, while reinforcement learning policies have been manipulated to produce suboptimal or dangerous actions during critical decision-making phases [15, 11]. Sequential perturbation attacks against quadrupedal locomotion systems have shown how adversaries can induce gait instabilities that compromise mobility, while supply chain attacks have been demonstrated against vision-language models used for robotic manipulation tasks [87, 110]. These studies indicated that while attack methodologies are well-documented across the surveyed papers, corresponding defense mechanisms appear in a few, creating a significant vulnerability gap. Moreover, these studies focus exclusively on isolated AI components without considering how AML attacks might propagate across integrated robotic architectures.
Middleware and Communication Security. Within the reviewed papers, ROS and data distribution service (DDS) security were among the most studied topics, with approximately one-third of the surveyed publications addressing both attack vectors and corresponding defense mechanisms [19, 82, 3]. Researchers have found that ROS and DDS are vulnerable to attacks like unauthorized command injection [96, 123], and some defenses such as intrusion and anomaly detection [94] have been proposed. However, the current research treats middleware as an isolated layer, failing to model how communication vulnerabilities interact with sensing, planning, and actuation subsystems in real-world deployments.
@@ 2.2 Recurring Architectural Themes
Our systematic analysis uncovered a distinct pattern: despite the apparent diversity of humanoid platforms and attack methodologies examined across papers, security concerns consistently clustered around seven distinct functional boundaries. During our content analysis, threats that initially appeared unrelated; consistently aligned with the same architectural layers. For example, sensor-spoofing attacks, whether targeting LiDAR systems [89, 14] or cameras [102, 93], invariably exploited vulnerabilities at the sensory interface between physical hardware and digital processing. These attacks succeeded by manipulating the boundary where analog sensor data transforms into digital representations used by higher-level algorithms. Similarly, AML attacks consistently targeted high-level reasoning components, exploiting the interface between data processing and decision-making logic [15, 11, 87]. Physical components, upon which all other system functions depend, can be compromised by hardware attacks. Most significantly, the observed patterns reflect real architectural boundaries in humanoid robots-not merely the areas researchers emphasized. Unlike traditional IT systems with arbitrary software boundaries, humanoids have natural layers that reflect their physical and digital components. These layers create consistent attack points where different subsystems connect, making them common targets across different platforms.
@@ 2.3 Critical Gaps
Our analysis shows that robotics security studies intensely focus on specific attack vectors or individual subsystems; only a few of the surveyed papers examined attacks spanning more than one subsystem or architectural layer. As a result, prior work often fails to capture realistic adversarial dynamics, where a local fault can propagate across the humanoid layers. We observe an asymmetric defense landscape. Human–robot interface (HRI) vulnerabilities are well documented [123, 120], but practical, validated mitigations are scarce. By contrast, several sensor-level attacks (e.g., LiDAR spoofing, IMU drift) [89, 14] already have concrete countermeasures (filtering, fusion checks, shielding, freshness tests) [109, 117]. The result is an imbalanced posture: upper layers that shape user intent are weakly protected, while lower layers have spot-fixes. Modern robotic systems need more measures like integrated defenses that work together across all system layers. This means binding human-robot interaction policies to authenticated, session-verified middleware; requiring sensor validation before any planning occurs; and implementing hard limits at the control level with automated recovery capabilities. The proposed model makes this integration measurable by identifying the specific layers and interfaces that must connect, while RISK-MAP quantifies whether a platform actually implements these connections and calculates the residual risk when any link fails.
@@ 2.4 Contributions of This Paper
    Novel seven layered model: Introduction of seven-layer security model tailored to humanoids that organizes 39 attack vectors and 35 defenses across natural system boundaries identified through our literature analysis.
@@
    Quantitative assessment method: We distill from the literature, a risk-weighted scoring system with statistical validation that enables systematic security evaluation and cross-platform benchmarking.
@@ 3.1 Architectural Analysis
The network-CPS duality. Our analysis uncovered that humanoids exist at the intersection of two established computational paradigms. Like networked systems, they use layered abstractions with defined interfaces, protocols, and modular data flows. Like CPSs, they couple computation to physical processes under real-time and safety constraints. This duality implies that an effective security architecture must treat interfaces as enforcement points for identity, authorization, integrity, confidentiality, and auditability and practically, use each interface to enforce who can communicate, what can be sent, whether it is within safe bounds, and recorded for audit.
@@
Therefore, we developed a seven-layer security architecture tailored to humanoids by refining the OSI model, in an iterative process. This process involved the following steps:
@@
    Defense Integration: Examining how the 35 defense mechanisms in the literature review can be systematically organized in the layered model.
@@
3.2 Seven-Layer Humanoid Security Model
@@ 4.1 Holistic Humanoid Threat Landscape
Real-time constraints make latency itself an attack vector. In enterprise IT, a few seconds of delay in anomaly detection may be tolerable; in robotics, locomotion and manipulation often run faster than 10 milliseconds, so the same delay can cause physical collapse, hardware damage, or injury. Adversaries can exploit this narrow operational window, for example, LiDAR spoofing (SP-A1) can destabilize motion within a control cycle, before a detector reacts. Security mechanisms that do not operate within these deadlines provide no prevention-they enable only post-incident forensic analysis.